# MAP: Unleashing Hybrid Mamba-Transformer Vision Backbone’s Potential with Masked Autoregressive Pretraining

### v1

@@ 4 Masked Autoregressive Pretraining for Hybrid Backbones
MAP Hybrid Mamba-Transformer Encoder. We designed a series of hybrid Mamba-Transformer vision backbones and compared their performance when trained from scratch. The results indicate that the hybrid approach using MMMTMMMT performs the best. When comparing Mamba-R* with MMMMMMTT, we found that adding a Transformer after Mamba enhances its long-context modeling capabilities, leading to improved performance. However, when comparing MMMMMMTT with TTMMMMMM, we observed that simply appending Transformers after Mamba does not fully leverage the architecture’s potential. This suggests that incorporating Transformers at the beginning is crucial for extracting sufficient local features. We believe that the MMMTMMMT approach effectively balances local feature extraction and contextual modeling enhancement, making it our default configuration.
@@
Figure 6: Different Hybrid Model Design. (d) achieves the best results and is set as default.
@@
Table 5: Hybrid Design of Mamba-Transformer backbone. All experiments are trained from scratch. Mamba-R* means 24 Mamba-R(Wang et al., 2024) Mamba layers plus 8 additional Mamba layers. DeiT* means 24 DeiT(Touvron et al., 2021) Transformer layers plus 8 additional Transformer layers. MMMMMMTT represents 24 Mamba layers followed by 8 Transformer layers. TTMMMMMM represents 8 Transformer layers followed by 24 Mamba layers. TMMMTMMM represents a unit consisting of 1 Transformer layer and 3 Mamba layers, repeated 8 times. MMMTMMMT represents a unit of 3 Mamba layers followed by 1 Transformer layer, repeated 8 times.
@@
MAP Transformer Decoder. To reconstruct the original image, we utilize a masked Transformer for signal recovery. Our decoder, while consistent with MAE, employs a distinct row-wise decoding strategy that allows autoregressive decoding of one row of tokens at a time, enhancing the network’s ability to capture local features and contextual relationships among regions. Experiments show that this method significantly outperforms the original AR, MAE, and local MAE decoding strategies. Notably, in the hybrid framework, local MAE performs comparably to standard MAE, emphasizing the significance of local feature learning. Our MAP method improves local feature modeling while leveraging autoregressive techniques to capture contextual relationships across regions, resulting in superior performance.
@@
Decoder Mask Autoregressive(AR) MAE local MAE MAP (ours)
@@
Reconstruction Target. Consistent with MAE, we reconstructed normalized original pixels as the target and employed MSE loss. Inspired by MAR(Li et al., 2024) to use reconstruction output as a conditional signal for diffusion models to improve generation quality, we explored whether pretraining with diffusion loss could enhance performance. However, this approach did not yield significant improvements. This may be due to the decoder’s increased capacity negatively impacting the encoder’s pretraining effectiveness, suggesting that the quality of reconstructed images is not directly linked to encoder pretraining success.
Reconstruction Target From Scratch Diffusion Loss MSE Loss (ours)
@@
Table 7: Reconstruction Target. Results indicate that the quality of the reconstructed image is not directly related to the pretraining effectiveness.
@@
ResNet-50 (He et al., 2016) 2242 25M 2388 6.6G 76.2
ResNet-152 (He et al., 2016) 2242 60M 1169 12.5G 78.3
EfficientNet-B3 (Tan & Le, 2019) 3002 12M 496 19.7G 81.6
ConvNeXt-T (Liu et al., 2022b) 2242 29M 701 8.3G 82.1
ConvNeXt-S (Liu et al., 2022b) 2242 50M 444 13.1G 83.1
ConvNeXt-B (Liu et al., 2022b) 2242 89M 334 17.9G 83.8
@@
Vim-T (Zhu et al., 2024a) 2242 7M 1165 4.8G 76.1
Vim-S (Zhu et al., 2024a) 2242 26M 612 9.4G 80.5
MambaR-T (Wang et al., 2024) 2242 9M 1160 5.1G 77.4
MambaR-S (Wang et al., 2024) 2242 28M 608 9.9G 81.1
MambaR-B (Wang et al., 2024) 2242 99M 315 20.3G 82.9
MambaR-L (Wang et al., 2024) 2242 341M 92 55.5G 83.2
@@
ARM-B (Mamba+AR) (Ren et al., 2024) 2242 85M 325 19.7G 83.2
ARM-L (Mamba+AR) (Ren et al., 2024) 2242 297M 111 53.1G 84.5
@@
VMamba-T (Liu et al., 2024) 2242 31M 464 7.6G 82.5
VMamba-S (Liu et al., 2024) 2242 50M 313 27.6G 83.6
VMamba-B (Liu et al., 2024) 2242 89M 246 37.1G 83.9
@@
MambaVision-T (Hatamizadeh & Kautz, 2024) 2242 35M 1349 10.7G 82.7
MambaVision-S (Hatamizadeh & Kautz, 2024) 2242 51M 1058 36.6G 83.3
MambaVision-B (Hatamizadeh & Kautz, 2024) 2242 97M 826 50.8G 84.2
MambaVision-L  (Hatamizadeh & Kautz, 2024) 2242 241M 229 78.6G 85.3
@@
HybridMH-T 2242 12M 910 7.6G 77.7
HybridMH-S 2242 37M 512 14.6G 81.3
HybridMH-B 2242 128M 244 30.0G 83.1
@@
HybridMH-T + MAP 2242 12M 910 7.6G 78.6
HybridMH-S + MAP 2242 37M 512 14.6G 82.5
HybridMH-B + MAE 2242 128M 244 30.0G 83.9
HybridMH-B + AR 2242 128M 244 30.0G 83.8
HybridMH-B + CL 2242 128M 244 30.0G 83.1
HybridMH-B + MAP 2242 128M 244 30.0G 84.9
@@
Table 8: ImageNet-1k classification results. The throughput is computed on an A100 GPU. The memory overhead is measured with a batch size of 128 on single GPU. Our results are highlighted in blue. Our proposed MAP method significantly improves the performance of the hybrid Mamba-Transformer backbones. Additionally, we verified that our MAP method also significantly improves the performance of both the pure Mamba framework and the pure Transformer backbone. Our MAP method also significantly outperforms MAE, AR, and CL pretraining on hybrid networks.
Hybrid Ratio 3M1T 3M1T+MAP 1M3T 1M3T+MAP 2M2T 2M2T+MAP
Accuracy 83.1 84.9 83.3 85.1 83.5 84.9
Table 9: Results on Different Hybrid Ratio. 3M1T denotes a ratio of 3:1 for Mamba and Transformer, while 3M1T+MAP indicates that it undergoes MAP pretraining first. The results reveal minimal performance differences among the various hybrid ratios after pertaining. Considering computational efficiency and memory savings, we use the 3:1 hybrid ratio as our default setting.
Method PT #P $\downarrow$ #F $\downarrow$ ScanObjectNN ModelNet40
OBJ_BG $\uparrow$ OBJ_ONLY $\uparrow$ PB_T50_RS $\uparrow$ 1k P $\uparrow$
@@
PointNet(Qi et al., 2017a) $\times$ 3.5 0.5 73.3 79.2 68.0 89.2
PointNet++(Qi et al., 2017b) $\times$ 1.5 1.7 82.3 84.3 77.9 90.7
DGCNN(Wang et al., 2019) $\times$ 1.8 2.4 82.8 86.2 78.1 92.9
PointCNN(Li et al., 2018) $\times$ 0.6 - 86.1 85.5 78.5 92.2
DRNet (Qiu et al., 2021) $\times$ - - - - 80.3 93.1
SimpleView(Goyal et al., 2021) $\times$ - - - - 80.5$\pm$0.3 93.9
GBNet(Qiu et al., 2022) $\times$ 8.8 - - - 81.0 93.8
PRA-Ne(Cheng et al., 2021) $\times$ - 2.3 - - 81.0 93.7
MVTN(Hamdi et al., 2021) $\times$ 11.2 43.7 92.6 92.3 82.8 93.8
PointMLP(Ma et al., 2022) $\times$ 12.6 31.4 - - 85.4$\pm$0.3 94.5
PointNeXt(Qian et al., 2022) $\times$ 1.4 3.6 - - 87.7$\pm$0.4 94.0
P2P-HorNet(Wang et al., 2022) $\checkmark$ - 34.6 - - 89.3 94.0
DeLA(Chen et al., 2023) $\times$ 5.3 1.5 - - 90.4 94.0
@@
Transformer $\times$ 22.1 4.8 79.86 80.55 77.24 91.4
PCT(Guo et al., 2021) $\times$ 2.9 2.3 - - - 93.2
PointMamba $\times$ 12.3 3.6 88.30 87.78 82.48 -
PCM(Zhang et al., 2024) $\times$ 34.2 45.0 - - 88.10$\pm$0.3 93.4$\pm$0.2
SPoTr(Park et al., 2023) $\times$ 1.7 10.8 - - 88.60 -
PointConT(Liu et al., 2023) $\times$ - - - - 90.30 93.5
Mamba3d w/o vot. $\times$ 16.9 3.9 92.94 92.08 91.81 93.4
Mamba3d w/ vot. $\times$ 16.9 3.9 94.49 92.43 92.64 94.1
HybridMT3D w/o vot. $\times$ 19.3 4.4 92.81 92.28 91.97 93.5
HybridMT3D w/ vot. $\times$ 19.3 4.4 94.50 92.58 92.66 94.3
@@
Transformer OcCo 22.1 4.8 84.85 85.54 78.79 92.1
Point-BERT IDPT 22.1+1.7 4.8 88.12 88.30 83.69 93.4
MaskPoint MaskPoint 22.1 4.8 89.30 88.10 84.30 93.8
PointMamba Point-MAE 12.3 3.6 90.71 88.47 84.87 -
Point-MAE IDPT 22.1+1.7 4.8 91.22 90.02 84.94 94.4
Point-M2AE Point-M2AE 15.3 3.6 91.22 88.81 86.43 94.0
Mamba3d w/o vot. Point-BERT 16.9 3.9 92.25 91.05 90.11 94.4
Point-MAE Point-MAE 22.1 4.8 90.02 88.29 85.18 93.8
Mamba3d w/o vot. Point-MAE 16.9 3.9 93.12 92.08 92.05 94.7
Mamba3d w/ vot. Point-MAE 16.9 3.9 95.18 94.15 93.05 95.4
Mamba3d w/o vot. MAP 16.9 3.9 93.62 92.75 92.65 95.1
Mamba3d w/ vot. MAP 16.9 3.9 95.64 94.87 93.76 95.6
HybridMT3D w/o vot. MAP 19.3 4.4 93.88 93.03 92.95 95.4
HybridMT3D w/ vot. MAP 19.3 4.4 95.84 94.97 93.87 95.9
Table 10: Results on 3D classification tasks. Our results are highlighted in blue.

### v2

@@ 4 Method
Preliminary.The focus of this paper is on studying how to pretrain hybrid Mamba-Transformer frameworks, rather than on designing the hybrid frameworks themselves. Therefore, we first determine a hybrid network as the default of our study. We tried a series of hybrid Mamba-Transformer vision backbones and compared their performance when trained from scratch. The results indicate that the hybrid approach using MMMTMMMT performs the best. When comparing Mamba-R* with MMMMMMTT, we found that adding a Transformer after Mamba enhances its long-context modeling capabilities, leading to improved performance. However, when comparing MMMMMMTT with TTMMMMMM, we observed that simply appending Transformers after Mamba does not fully leverage the architecture’s potential. This suggests that incorporating Transformers at the beginning is crucial for extracting sufficient local features. We believe that the MMMTMMMT approach effectively balances local feature extraction and contextual modeling enhancement, making it our default configuration. We adopt the network shown in Figure 4(d) as our default hybrid network and refer to it as HybridNet. It is worth noting that our method can also be applied to other Mamba-Transformer hybrid frameworks, but the design of the hybrid frameworks themselves is not within the scope of this paper.
@@
Figure 4: Different Hybrid Model Design. (d) achieves the best results and is set as default and refer to it as HybridNet.
@@
Table 4: Hybrid Design of Mamba-Transformer backbone. All experiments are trained from scratch. Mamba-R* means 24 Mamba-R[46] Mamba layers plus 8 additional Mamba layers. DeiT* means 24 DeiT[43] Transformer layers plus 8 additional Transformer layers. MMMMMMTT represents 24 Mamba layers followed by 8 Transformer layers. TTMMMMMM represents 8 Transformer layers followed by 24 Mamba layers. TMMMTMMM represents a unit consisting of 1 Transformer layer and 3 Mamba layers, repeated 8 times. MMMTMMMT represents a unit of 3 Mamba layers followed by 1 Transformer layer, repeated 8 times.
Overview. Our proposed MAP is a pre-training strategy for the hybrid Mamba-Transformer framework that can significantly improve the performance of the hybrid architecture compared to traditional MAE and AR pre-training. Our MAP pre-training is to perform random masking on a given image and then region-wise autoregressively reconstruct its original image. Here, the tokens of each row of the image are predicted at the same time, and the tokens between rows are predicted by autoregression.
Notably, we choose each row as a sub-region here because, as mentioned in the previous chapter, the AR order of Mamba should match the scanning order. The default scanning order for most Mamba implementations is row-first. To ensure a fairer comparison with existing methods, we treat each row as a prediction unit. After changing the scanning order of Mamba, more sophisticated clustering strategies are theoretically expected to yield better results.
As shown in Figure 2(c), given an image, HybridNet first maps the randomly masked image into the feature space and then uses a Transformer Decoder to decode the original image row-wise. Consider an image $\mathbf{I}$ which is divided into rows:
$\mathbf{I}=\{\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{M}\}$ (1)
Each row $\mathbf{r}_{i}$ consists of tokens:
$\mathbf{r}_{i}=\{\mathbf{x}_{i1},\mathbf{x}_{i2},\ldots,\mathbf{x}_{iN}\}$ (2)
Select a subset of tokens to mask within each row. Let $\mathbf{M}_{i}\subset\{1,2,\ldots,N\}$ denote the indices of the masked tokens in row $\mathbf{r}_{i}$.
For a given row $\mathbf{r}_{i}$, predict all masked tokens simultaneously:
$p(\mathbf{x}_{ij}\mid\mathbf{x}_{i,j\notin\mathbf{M}_{i}},\mathbf{r}_{<i})$ (3)
where $\mathbf{r}_{<i}$ refers to all rows preceding row $i$.
The prediction of tokens in row $i$ depends on all previous rows and the visible tokens within the row. This can be expressed as:
$p(\mathbf{r}_{i}\mid\mathbf{r}_{<i})=\prod_{j=1}^{N}p(\mathbf{x}_{ij}\mid\mathbf{x}_{i,j\notin\mathbf{M}_{i}},\mathbf{r}_{<i})$ (4)
The overall loss function is the sum of the negative log-likelihoods of the predicted tokens:
$\mathcal{L}=-\sum_{i=1}^{M}\sum_{j\in\mathbf{M}_{i}}\log p(\mathbf{x}_{ij}\mid\mathbf{x}_{i,j\notin\mathbf{M}_{i}},\mathbf{r}_{<i})$ (5)
Our proposed MAP represents a general paradigm applicable to data across various domains, with 2D image data as an example. Our method can be easily extended to large language models and the fields of image video and point cloud video. Our method optimizes the synergy between Mamba and Transformer within a unified framework, allowing both models to fully leverage their strengths. In the Mamba-Transformer hybrid architecture, this approach effectively enhances the cooperation between the models, resulting in significant performance improvements.
For HybridNet, the masking mechanism of MAP can significantly enhance the local modeling capabilities of both the Mamba and Transformer layers. The region-wise autoregressive strategy of MAP can significantly improve Mamba’s ability to model context. Therefore, overall, MAP can bring about substantial performance improvements for HybridNet. For pure Mamba architecture, our method outperforms AR, due to MAP’s ability to effectively model local features and the associations between local regions. On the pure Transformer architecture, although MAP does not surpass the performance of MAE, it still achieves comparable results. This is because, while local MAE reduces the receptive field, the autoregressive modeling between local regions enhances the ability to capture local relationships, ensuring that MAP maintains strong performance on Transformers.
We will introduce the specific design components of the framework. The subsequent experiments in this section are conducted using the base-sized model on the ImageNet-1K dataset.
Masking. Since masking is also crucial for MAE pretraining of Transformers and AR pretraining of Mamba, masking is an important design aspect of MAP. We experimented with different masking strategies, including random, sequential, and diagonal masking. Our experiments show that random masking delivers the best results. We attribute this to the fact that sequential and diagonal masking can hinder the Transformer’s ability to establish contextual relationships. Random masking not only promotes bidirectional modeling for Transformers but also enhances Mamba’s generalization and representation capabilities in sequence modeling. The random masking strategy is actually a balance, as it retains the bidirectional modeling capability of Transformers while leveraging row-wise autoregressive pretraining to train Mamba. The raster and diagonal strategies fail to fully utilize the bidirectional modeling capability of Transformers and do not enhance Mamba’s ability to model local features.
Additionally, we explored the effects of different masking ratios and found that a 50% masking ratio yielded the best results. This conclusion aligns with intuition: while MAE performs optimally on Transformers with a 75% masking ratio, previous experiments showed that AR achieves the best results on Mamba with a 20% ratio. Therefore, a 50% ratio serves as a balanced number, leveraging the strengths of both paradigms. This ratio is significantly different from the conclusions drawn in MAE.
Strategy FS Random Sequential Diagonal
Accuracy 83.1 84.9 84.0 83.8
Ratio 0% 25% 50% 75%
Accuracy 83.3 84.5 84.9 84.2
Table 5: Random masking with a 50% ratio performs the best.
@@
MAP Transformer Decoder. To reconstruct the original image, we utilize a masked Transformer for signal recovery. The reason for using a Transformer decoder rather than a Mamba decoder is that the Transformer decoder can reconstruct region-wise based on the encoder’s features by applying a decoder mask. In contrast, the Mamba decoder, due to its unidirectional scanning nature, struggles to simultaneously reconstruct an entire local region. Our decoder employs a distinct row-wise decoding strategy that allows autoregressive decoding of one row of tokens at a time, enhancing the network’s ability to capture local features and contextual relationships among regions. Experiments show that this method significantly outperforms the original AR, MAE, and local MAE decoding strategies. Local MAR refers to limiting the receptive field of MAE within each row of the image. Notably, in the hybrid framework, local MAE performs comparably to standard MAE, emphasizing the significance of local feature learning. Our MAP method improves local feature modeling while leveraging autoregressive techniques to capture contextual relationships across regions, resulting in superior performance.
@@
Decoder Mask AR MAE local MAE ours
@@
Reconstruction Target. Consistent with MAE, we reconstructed normalized original pixels as the target and employed MSE loss. Inspired by MAR[20] to use reconstruction output as a conditional signal for diffusion models to improve generation quality, we explored whether pretraining with diffusion loss could enhance performance. However, this approach did not yield significant improvements, suggesting that the quality of reconstructed images is not directly linked to encoder pretraining success.
Loss FS Diffusion Loss MSE Loss (ours)
@@
Table 7: Reconstruction Target.
5 Experiments
5.1 2D Classification on ImageNet-1k.
Settings. We pretrained on the training set of the ImageNet-1K[7] dataset and then fine-tuned on its classification task. We report the top-1 validation accuracy of a single 224x224 crop, and in some settings, we also report the results for a 384x384 crop. During the pretraining phase, we applied a random masking strategy with a 50% masking ratio, using only random cropping as the data augmentation strategy. We utilized AdamW as the optimizer and trained for 1600 epochs across all settings. In the fine-tuning phase, we directly fine-tune for 400 epochs and report the results.
@@ 5 Experiments
ResNet-50 [17] 2242 25M 2388 6.6G 76.2
ResNet-152 [17] 2242 60M 1169 12.5G 78.3
EfficientNet-B3 [42] 3002 12M 496 19.7G 81.6
ConvNeXt-T [31] 2242 29M 701 8.3G 82.1
ConvNeXt-S [31] 2242 50M 444 13.1G 83.1
ConvNeXt-B [31] 2242 89M 334 17.9G 83.8
@@
Vim-T [63] 2242 7M 1165 4.8G 76.1
Vim-S [63] 2242 26M 612 9.4G 80.5
MambaR-T [46] 2242 9M 1160 5.1G 77.4
MambaR-S [46] 2242 28M 608 9.9G 81.1
MambaR-B [46] 2242 99M 315 20.3G 82.9
MambaR-L [46] 2242 341M 92 55.5G 83.2
@@
ARM-B (Mamba+AR) [41] 2242 85M 325 19.7G 83.2
ARM-L (Mamba+AR) [41] 2242 297M 111 53.1G 84.5
MambaR-B+MAE 2242 99M 315 20.3G 83.1
MambaR-B+AR 2242 99M 315 20.3G 83.7
@@
VMamba-T [29] 2242 31M 464 7.6G 82.5
VMamba-S [29] 2242 50M 313 27.6G 83.6
VMamba-B [29] 2242 89M 246 37.1G 83.9
@@
MambaVision-T [16] 2242 35M 1349 10.7G 82.7
MambaVision-S [16] 2242 51M 1058 36.6G 83.3
MambaVision-B [16] 2242 97M 826 50.8G 84.2
MambaVision-L  [16] 2242 241M 229 78.6G 85.3
MambaVision-B+MAP [16] 2242 97M 826 50.8G 84.9
MambaVision-L+MAP  [16] 2242 241M 229 78.6G 86.4
@@
HybridNet-T 2242 12M 910 7.6G 77.7
HybridNet-S 2242 37M 512 14.6G 81.3
HybridNet-B 2242 128M 244 30.0G 83.1
@@
HybridNet-T + MAP 2242 12M 910 7.6G 78.6
HybridNet-S + MAP 2242 37M 512 14.6G 82.5
HybridNet-B + MAE 2242 128M 244 30.0G 83.9
HybridNet-B + AR 2242 128M 244 30.0G 83.8
HybridNet-B + CL 2242 128M 244 30.0G 83.1
HybridNet-B + MAP 2242 128M 244 30.0G 84.9
@@
Table 8: ImageNet-1k classification results. The throughput is computed on an A100 GPU. The memory overhead is measured with a batch size of 128. Our results are highlighted in blue.
Results. Results are shown in Table 8. The results indicate that the hybrid framework achieves a balance between performance and computational overhead. The focus of this paper is not on designing the hybrid network but on investigating its pretraining methods. Our proposed pretraining method MAP significantly enhances the performance of the hybrid Mamba-Transformer framework. Comparing the performance of MAP, MAE, AR, and CL on HybridNet-B, it is evident that MAP significantly improves performance. This indicates that compared to MAE, the local MAE in MAP, which enhances local modeling and adopts an autoregressive approach to strengthen the associations between local regions, is crucial for hybrid networks. Compared to AR, the masking mechanism and the local autoregressive strategy in MAP contribute to performance gains. Additionally, we verified that our MAP method can improve the performance of the pure Mamba framework and the pure Transformer backbone. By comparing AR, MAE, and MAP pretraining results on MambaR-B, we can see that MAE offers limited improvement for pure Mamba. AR, benefiting from its autoregressive modeling of context, shows improved performance. Compared to AR, the masking mechanism and local autoregressive decoding strategy in MAP enhance Mamba’s ability to model local features, further improving performance. By comparing the performance of AR, MAE, and MAP on ViT, we find that on ViT-B, MAP can achieve the same results as MAE. This is because, although local MAE reduces the receptive field compared to the original MAE, region-wise autoregressive modeling not only enhances local feature modeling but also strengthens the modeling of correlations between local regions, preventing a decline in performance. When comparing MAE and MAP on ViT-L, the advantage of region-wise autoregressive modeling in MAP becomes apparent in larger-scale models, where it outperforms MAE. This also suggests that autoregressive modeling has advantages in larger models, which is consistent with the scaling law phenomenon observed in LLMs. In addition, by comparing the results of HybridNet-B/L at 224 and 384 resolutions, we can see that at the longer sequence length of 384 resolution, our method further achieved improvements of 0.6 and 0.8, respectively. This indicates that the longer context modeling allowed by the Mamba layer can indeed lead to performance gains. Comparing the results of MambaVision-B/L before and after MAP pretraining shows that our MAP method not only benefits the HybridNet used as an example but also brings significant improvements to other existing hybrid Mamba-Transformer network frameworks.
5.2 Transfer Learning Experiments
Semantic Segmentation We conduct experiments for semantic segmentation on the ADE20K [61] and use UperNet [53] as the segmentation framework followed by Vim. All settings are consistent with Vim. When comparing HybridNet-S with Vim-S, we can see that introducing a Transformer significantly enhances the performance of dense prediction tasks. This highlights the value of hybrid networks, which maintain satisfactory performance when introducing longer contexts at 512x512 resolution. Comparing the results of HybridNet-S before and after MAP pretraining, we can see that our pretraining method also brings significant performance improvements in segmentation tasks.
Object Detection and Instance Segmentation We conduct experiments for object detection and instance segmentation on the COCO 2017 dataset [25] and use ViTDet [53] as the basic framework followed by Vim. All settings are consistent with Vim. The results indicate that, while simply inserting a Transformer into Vim does not bring about significant improvements, the performance of HybridNet-Ti, when fine-tuned after initialization with MAP pretraining, shows a notable improvement.
Method Backbone
image
size
#param.
$val$
mIoU
DeepLab v3+ ResNet-101 $512^{2}$ 63M 44.1
UperNet ResNet-50 $512^{2}$ 67M 41.2
UperNet ResNet-101 $512^{2}$ 86M 44.9
UperNet DeiT-Ti $512^{2}$ 11M 39.2
UperNet DeiT-S $512^{2}$ 43M 44.0
UperNet Vim-Ti $512^{2}$ 13M 41.0
UperNet Vim-S $512^{2}$ 46M 44.9
UperNet HybridNet-S $512^{2}$ 69M 45.6
UperNet HybridNet-S+MAP $512^{2}$ 69M 46.9
Table 9: Results of semantic segmentation on the ADE20K $val$ set.
[…]
DeiT-Ti 44.4 38.1
Vim-Ti 45.7 39.2
HybridNet-Ti 45.9 39.2
HybridNet-Ti+MAP 46.4 39.8
Table 10: Results of object detection and instance segmentation on the COCO $val$ set using Cascade Mask R-CNN [2] framework.
Method PT #P $\downarrow$ #F $\downarrow$ ModelNet40
@@ 5.2 Transfer Learning Experiments
PointNet[36] $\times$ 3.5 0.5 89.2
PointNet++[37] $\times$ 1.5 1.7 90.7
DGCNN[50] $\times$ 1.8 2.4 92.9
PointCNN[21] $\times$ 0.6 - 92.2
DRNet [39] $\times$ - - 93.1
SimpleView[10] $\times$ - - 93.9
GBNet[40] $\times$ 8.8 - 93.8
PRA-Ne[6] $\times$ - 2.3 93.7
MVTN[13] $\times$ 11.2 43.7 93.8
PointMLP[32] $\times$ 12.6 31.4 94.5
PointNeXt[38] $\times$ 1.4 3.6 94.0
P2P-HorNet[51] $\checkmark$ - 34.6 94.0
DeLA[4] $\times$ 5.3 1.5 94.0
@@
Transformer $\times$ 22.1 4.8 91.4
PCT[12] $\times$ 2.9 2.3 93.2
PointMamba[23] $\times$ 12.3 3.6 -
PCM[60] $\times$ 34.2 45.0 93.4$\pm$0.2
SPoTr[35] $\times$ 1.7 10.8 -
PointConT[28] $\times$ - - 93.5
Mamba3d w/o vot.[15] $\times$ 16.9 3.9 93.4
Mamba3d w/ vot.[15] $\times$ 16.9 3.9 94.1
HybridNet w/o vot. $\times$ 19.3 4.4 93.5
HybridNet w/ vot. $\times$ 19.3 4.4 94.3
@@
Transformer OcCo[48] 22.1 4.8 92.1
Point-BERT IDPT[58] 22.1+1.7 4.8 93.4
MaskPoint MaskPoint[27] 22.1 4.8 93.8
PointMamba Point-MAE[33] 12.3 3.6 -
Point-MAE IDPT[58] 22.1+1.7 4.8 94.4
Point-M2AE Point-M2AE[59] 15.3 3.6 94.0
Mamba3d w/o vot. Point-BERT[57] 16.9 3.9 94.4
Point-MAE Point-MAE[33] 22.1 4.8 93.8
Mamba3d w/o vot. Point-MAE[33] 16.9 3.9 94.7
Mamba3d w/ vot. Point-MAE[33] 16.9 3.9 95.4
Mamba3d w/o vot. MAP 16.9 3.9 95.1
Mamba3d w/ vot. MAP 16.9 3.9 95.6
HybridNet w/o vot. MAP 19.3 4.4 95.4
HybridNet w/ vot. MAP 19.3 4.4 95.9
Table 11: Results on 3D ModelNet classification tasks.
5.3 3D Classification Experiments
Settings. To verify that our method can be applied to other data formats, we transferred the MAP pretraining to 3D data for experimentation. We pretrained using the ShapeNet[3] dataset, employing random rotation and translation scaling as data augmentation techniques. Each point cloud consists of 1024 points and is divided into 64 patches, with each patch containing 32 points. Since point clouds are unordered, the concept of rows does not apply here; instead, we randomly generate 32 patches each time and complete the reconstruction process in an autoregressive manner. Similar to Mamba3D[15], we did not adopt any special sorting strategies but ensured that the order of pretraining matches that of the actual Mamba scans. We conducted pretraining on both the hybrid framework and the original Mamba3D to validate their performance advantages. During pretraining and downstream fine-tuning, we employed the AdamW optimizer with a cosine decay strategy for 300 epochs.
Results.The experiments demonstrate that our method also can improve the performance on 3D tasks. This suggests that our approach can be easily adapted to other domains. Comparing HybridNet and Mamba3d at the backbone network level, it is evident that the performance improvement solely from inserting a Transformer is limited. By comparing the results of HybridNet before and after MAP pretraining, we can see that our MAP pretraining brings about a significant improvement. This further validates the effectiveness of our proposed pretraining method. More experiments on 3D datasets can be found in the Supplementary Material.
6 Conclusion
In this paper, we propose Masked Autoregressive Pretraining, a pretraining framework suitable for hybrid Mamba-Transformer models. It integrates the advantages of MAE and AR, retaining their core designs to model local features while strengthening the modeling of contextual correlations. Extensive experiments have validated the effectiveness of our method and the rationality of its component design.
Supplementary Material
Additional Experiments on 3D Tasks
To verify that our method not only performs well on 2D tasks but is also effective for 3D tasks, we conducted experiments on additional 3D tasks. In addition to the ModelNet40[52] classification experiment, we also tested on the more challenging ScanObjectNN[44] dataset. Unlike ModelNet40, which is a synthetic dataset, ScanObjectNN is a real-world object dataset and is commonly evaluated under three settings: OBJ_BG, OBJ_ONLY, and PB_T50_RS. Among these, the PB_T50_RS setting is the most challenging. Comparing the results of HybridNet and Mamba3D under the Supervised Learning Only setting reveals that HybridNet performs only slightly better than Mamba3D. However, both HybridNet and Mamba3D achieve significant performance improvements after MAP pretraining. This further validates that the MAP pretraining strategy is not only effective for hybrid frameworks but also enhances the pure Mamba framework. Comparing the results of Mamba3D under Point-BERT, Point-MAE, and MAP, it is evident that MAP demonstrates a significant performance advantage. This proves that even within the pure Mamba framework, MAP can surpass the performance of BERT-style and MAE-style pertaining. We also conducted experiments on the few-shot learning task of ModelNet40 to validate the effectiveness of MAP. After MAP pretraining, both HybridNet and Mamba3D achieved significant performance improvements. On the more fine-grained task of ShapeNetPart[55] part segmentation, we also demonstrated that MAP can bring significant performance improvements to both hybrid frameworks and the pure Mamba framework.